The perimeter of a rectangle is six times the width. If the length were increased by 4 inches and the width were increased by 7 inches, then the perimeter would be 160 inches. What are the length and the width?
Answer by mananth(16946)

let the width be x
the perimeter = 6x
sum of the lengths = 6x-2x = 4x
each lengh of rectangle = 2x
2*(2x+4)+2*(x+7)=160
4x+8 + 2x+14 =160
^x+22 =160
6x=138
x= 138/3
x= 23 the width
length = 2x = 46




(